OpenStudy (anonymous):

16 over 12 - 14 0ver x+1

OpenStudy (anonymous):

\[\frac{16}{12}-\frac{14}{x+1}\] we need to find a common denominator

OpenStudy (anonymous):

first we can reduce 16/12=4/3

OpenStudy (anonymous):

someone needs to get some sleep! is your goal to stay a master and use an alias forever?

OpenStudy (anonymous):

\[\frac{4}{3}-\frac{14}{x+1}=\frac{4(x+1)-14*3}{3(x+1)}=\frac{4x+4-42}{3x+3}=\frac{4x-38}{3x+3}\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i don't know if i should be a guru because of that horrible movie

OpenStudy (anonymous):

i have to say i like "master" better. and sensei is just kitsch

OpenStudy (anonymous):

what movie?

OpenStudy (anonymous):

omg who is that guy who was austin powers he played in a horible movie called the guru or guru

OpenStudy (anonymous):

oh lord. you must have been on of the 37 people who went to see the movie.

OpenStudy (anonymous):

but i know what you are talking about. i want to be baba au rhum das cake for my next level.

OpenStudy (anonymous):

no we illegally download it

OpenStudy (anonymous):

ok one of the 37 people who watched it.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

after level 100 i think we should get to pick our own honorific

OpenStudy (anonymous):

satellite i miss teaming up on people it seems we don't do it as much i have a new satellite now we do team ups

OpenStudy (anonymous):

new satellite! no fair

OpenStudy (anonymous):

his name is joe

OpenStudy (anonymous):

but i have somewhat less time now. maybe tonight we can team up again on some poor unsuspecting soul

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

if i'm not asleep i cant get back to my right sleep schedule

OpenStudy (anonymous):

\[\color{green}{\text{jealous}}\]

OpenStudy (anonymous):

off to the market for me. don't stay up all day

OpenStudy (anonymous):

soo the answer is ? lol

OpenStudy (anonymous):

i think it was like (4x-38)/(3x+3) if i remember right

OpenStudy (anonymous):

this was a subtraction not an equality yes?

OpenStudy (anonymous):

oh her others one have been equalitys havent they?

OpenStudy (anonymous):

answer choices are 9.5 2 4 10.5

OpenStudy (anonymous):

you have 16/12-14/(x+1) i suppose you did not mean this then

OpenStudy (anonymous):

does - suppose to an equal sign?

OpenStudy (anonymous):

no, thank you though :)

OpenStudy (anonymous):

if you have 4/3=14/(x+1) cross multiply 4(x+1)=42 (x+1)=42/4 x+1=10.5 x=10.5-1=9.5 but u said it is suppose to be a minus sign?